find the average annual growth rate for each of the firms

Firm 2006 2007 2008 2009 2010 2011
Loewen 1 1.07 1.3 1.21 1.27 1.35
Morse 1.12 0.8 0.8 1.2 1.35 1.35
Huddleston 1.75 2.5 3.6 3.75 3.8 4.25
Meyer 2.25 2.5 2.1 2.74 2.85 2.96
Answer #1

HI

For Average annual growth rate at first each year growth rate need to be calculated:

Below is the excel formula and excel values for average annual growth rate for these firms:

x ✓ fox A C D E F G H I 2006 Average annual growth rate Firm Loewen growth rate 2007 1.07 =E3/D3-1 2008 1.3 =F3/E3-1 2009 1.2

Values are below:

Firm 2006 2007 2008 2009 2010 2011 Average annual growth rate
Loewen 1 1.07 1.3 1.21 1.27 1.35
growth rate 7.00% 21.50% -6.92% 4.96% 6.30% 6.57%
Morse 1.12 0.8 0.8 1.2 1.35 1.35
growth rate ###### 0.00% 50.00% 12.50% 0.00% 6.79%
Huddleston 1.75 2.5 3.6 3.75 3.8 4.25
growth rate 42.86% 44.00% 4.17% 1.33% 11.84% 20.84%
Meyer 2.25 2.5 2.1 2.74 2.85 2.96
growth rate 11.11% -16.00% 30.48% 4.01% 3.86% 6.69%

Thanks